10.Find f(-1) for the equation \\(y=2xy^{2}-4x+1)\\
a.7
b.-7
c.-1
d.1
1
Expert's answer
2019-09-23T05:32:25-0400

Let у = 2x2 - 4x + 1

f (-1)?

у = 2 * (-1) 2 - 4 * (-1) + 1

у = 2 + 4 + 1 = 7

Answer: а. 7.


Let "y=2xy^2-4x+1," then "y=2*(-1)y^2-4*(-1)+1, \, 2y^2+y-5=0,"

"D=1-4*2*(-5)=41,\\ y_1=(-1-\sqrt{41})/(2*2),"

"y_2=(-1+\sqrt{41})/(2*2)".

Answer:

"f(-1)=(-1-\sqrt{41})/(2*2),\\ or \, f(-1)=(-1+\sqrt{41})/(2*2)."